8.-12 points SerCP11 15.3.P.020 My Notes Ask Your Teacher An electron is accelerated by a constant electric field of magnitude 320 N/C (a) Find the acceleration of the electron. (b) Use the equations of motion with constant acceleration to find the electrons speed after 1.20 x 10-8 s, assuming it starts from rest. m/s Need Help?Read ItWatch it
Answer #1

Solution)

Given, E=320 N/C

Part a) We know, Acceleration, A= Ee/m =320*(1.6*10^-19)/(9.1*10^-31) =5.62*10^13 m/s^2

======

Part b) Given, u=0

T=1.2*10^-8 secs

We know, v=u+at=(5.62*10^13)*(1.2*10^-8)=6.75*10^5 m/s (Ans)

======